OpenStudy (anonymous):

If the side lengths of a cube are 16 feet, what is the correct way to write the expression to represent the volume of the cube in exponential form? 16 ⋅ 3 16 ⋅ 16 ⋅ 16 16^3 3^16

OpenStudy (anonymous):

@King.Void. lol *gag* math but can u get ur freinds on this q asap

OpenStudy (anonymous):

@kelre713

OpenStudy (anonymous):

@jane11509

OpenStudy (king.void.):

@confluxepic

OpenStudy (confluxepic):

Yes. @King.Void.

OpenStudy (anonymous):

he wants u to help me

OpenStudy (anonymous):

fast plz

OpenStudy (anonymous):

or at least explay a example of expotentiol

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

@confluxepic

OpenStudy (confluxepic):

Volume of a cube is (x)(x)(x).

OpenStudy (confluxepic):

\[V=16 \times 16 \times 16\]

OpenStudy (anonymous):

k

OpenStudy (confluxepic):

Now what do you think the answer is. @pinkiepieinthetardis

OpenStudy (confluxepic):

There are three of the same.

OpenStudy (anonymous):

b

OpenStudy (confluxepic):

No. It says in exponential form. So if I have \(2 \times 2\) then it would be \(2^{2}\) because two is being multiplied twice.

OpenStudy (anonymous):

oh so its 16^3

OpenStudy (confluxepic):

If I had \(2 \times 2 \times 2 \times 2 \times 2\) it would be \(2^{5}\)

OpenStudy (confluxepic):

Yes. You are correct. It is C.

OpenStudy (anonymous):

thx

OpenStudy (confluxepic):

You're welcome.